Sodium benzoate (NaC7H5O2) is sometimes used as a preservative in foods. Benzoic acid is the conjugate acid of the benzoate ion, which has a Ka value of 6.3 x 10-5.  

  1. Calculate the Kb value for the benzoate ion. (1 mark)
  2. Use the Kb value for the benzoate ion from part a, calculate the pH of a 0.215 M solution of the sodium benzoate.  (Ignore x method, 4 marks)
Answer #1

Kb is calculated:

Kb = Kw / Ka = 10 ^ -14 / 6.3E-5 = 1.60E-10

You have the reaction with water:

C7H5O2- + H2O = HC7H5O2 + OH-

We have the expression of Kb:

Kb = [HC7H5O2] * [OH-] / [C7H5O2-]

1.60E-10 = X ^ 2 / 0.215

Clears:

X = [OH-] = 5.90E-6 M

The pOH and pH are calculated:

pOH = - log [OH-] = - log 5.90E-6 = 5.23

pH = 14 - pOH = 14 - 5.23 = 8.77

If you liked the answer, please rate it positively, you would help me a lot, thanks.